How do I find the values of more than one data in a set given the values of harmonic geometric and arithmetic mean i.e


In a study a set of observations were as follows:x,16,20b,4,25,c.Given that the geometric mean is 8.433 the arithmetic mean 11.25 and the harmonic mean 6.026 compute the values x y z

1
Expert's answer
2021-11-16T18:43:40-0500
"x,20y, 4, 25, z""GM=\sqrt[6]{x(16)(20y)(4)(25)(z)}=8.433""AM=\dfrac{x+16+20y+4+25+z}{6}=11.25""HM=\dfrac{6}{\dfrac{1}{x}+\dfrac{1}{16}+\dfrac{1}{20y}+\dfrac{1}{4}+\dfrac{1}{25}+\dfrac{1}{z}}=6.026""x(20y)z=224.787739""x+20y+z=22.5""\dfrac{1}{x}+\dfrac{1}{20y}+\dfrac{1}{z}=0.643185""x+20y+z=22.5""x(20y)+x(z)+20y(z)=144.580101"




"x(20y)z=224.787739"

Vieta's Theorem



"t^3-22.5t^2+144.580101t-224.787739=0"

Solve graphically


"t_1=2.285, t_2=8.166, t_3=12.049"

Answer


"x=2.285, y=0.4083, z=12.049"

Or



"x=2.285, y=0.60245, z=8.166"

Or



"x=8.166, y=0.11425, z=12.049"

Or



"x=8.166, y=0.60245, z=2.285"

Or



"x=12.049, y=0.11425, z=8.166"

Or



"x=12.049, y=0.4083, z=2.285"